### v1

# One-Shot Method for Computing Generalized Winding Numbers

Abstract
The generalized winding number is an essential part of the geometry processing toolkit, allowing to quantify how much a given point is inside a surface, often represented by a mesh or a point cloud, even when the surface is open, noisy, or non-manifold. Parameterized surfaces, which often contain intentional and unintentional gaps and imprecisions, would also benefit from a generalized winding number. Standard methods to compute it, however, rely on a surface integral, challenging to compute without surface discretization, leading to loss of precision characteristic of parametric surfaces.
We propose an alternative method to compute a generalized winding number, based only on the surface boundary and the intersections of a single ray with the surface. For parametric surfaces, we show that all the necessary operations can be done via a Sum-of-Squares (SOS) formulation, thus computing generalized winding numbers without surface discretization with machine precision. We show that by discretizing only the boundary of the surface, this becomes an efficient method.
We demonstrate an application of our method to the problem of computing a generalized winding number of a surface represented by a curve network, where each curve loop is surfaced via Laplace equation. We use the Boundary Element Method to express the solution as a parametric surface, allowing us to apply our method without meshing the surfaces. As a bonus, we also demonstrate that for meshes with many triangles and a simple boundary, our method is faster than the hierarchical evaluation of the generalized winding number while still being precise.
We validate our algorithms theoretically, numerically, and by demonstrating a gallery of results on a variety of parametric surfaces and meshes, as well uses in a variety of applications, including voxelizations and boolean operations.
Figure 1: We propose a novel one-shot method to compute generalized winding numbers that does not require discretization of the surface. Our method only uses the boundary of the surface and a single ray-surface intersection test. For parametric surfaces (a, Coons patches), we thus can compute generalized winding number with machine precision without meshing or sampling the surface. Similarly, we compute generalized winding numbers for a curve network surfaced with minimal surfaces to high precision (b). Finally, our method can be also applied to meshes (c), where for some meshes it can improve the performance over standard methods while staying exact. Piggy Bank mesh with minor modifications by belch.
1 Introduction
For meshes and point clouds, even representing open, noisy, and non-manifold geometry, the notion of generalized winding number captures how much a point is inside the surface Jacobson et al. (2013); Barill et al. (2018). Generalized winding numbers are used in a variety of applications, including tetrahedral meshing Hu et al. (2018), surface reconstruction Xu et al. (2023), and medical imaging Becciu et al. (2011), to name a few.
The standard methods of computing generalized winding numbers calculate the projected area of the surface onto a unit sphere. For parametric surfaces, which often contain gaps (Fig. 2a,b) and thus also may benefit from generalized winding number, computing such projected area requires surface discretization. This discretization may defeat the purpose of having a high-precision parametric surface, as it can lead to a significant loss of precision (Fig. 2c,d).
We propose an alternative method to compute generalized winding number addressing this challenge. Our main insight is that for any surface, with or without boundary, computing a generalized winding number at a point requires the surface boundary only and intersections of the surface with a single ray. We furthermore show that in typical scenarios, when processing multiple query points, one can use less than one ray per query point on average, significantly reducing time complexity.
Our method is generic and can be applied to various geometry representations. For parametric surfaces, our method allows us to compute the winding number to machine precision, bypassing surface discretization. We show that the key components of our method, including the ray-surface intersection, can be done exactly via a convex Sum-of-Squares formulation Marschner et al. (2021). In practice, if both precision and efficiency are important, we can discretize only the boundary of the parametric shape, and still get a high-precision result while being memory efficient and fast.
We demonstrate an application of our method to piecewise minimal surfaces on 3D curve networks, collections of 3D curves forming one or multiple loops (Fig. 1), that are easy to create in virtual or augmented reality (VR/AR) contexts Google (2023), and often used for quick modeling, ideation, and visualization Gryaditskaya et al. (2020) (Fig. 16). We apply our method to compute a winding number of piecewise minimal surfaces by expressing them in a parametric form via Boundary Element Method (BEM). As before, our method allows us to avoid discretizing these surfaces.
For meshes with few boundary edges, in typical application scenarios such as voxelization, our method can be significantly faster than the hierarchical evaluation of the winding number Jacobson et al. (2013). While an approximation can be computed more efficiently Barill et al. (2018), our computation is exact.
We showcase our method on a wide variety of standard tasks for generalized winding numbers for all these representations (Sec. 7), including parametric Coons patches and Bézier triangles. We extensively validate our method by both formally showing it is equal to the generalized winding number and validating its numerical precision and performance in numerous experiments, as compared to Jacobson et al. (2013); Barill et al. (2018) (Sec. 7).
Our contributions are:
  • demonstrating that a computation of generalized winding number requires boundary-only operations and finding intersections of a surface with a single ray and that the same ray can be reused for multiple query points,
  • a new method of computing generalized winding number based on this observation, enabling exact computations for parametric surfaces, fast computations for some meshes, and computations for minimal surfaces requiring no meshing.
Figure 2: Parametric surfaces often contain gaps and overlaps, either intentional for tight alignment after manufacturing (a) or unintentional due to modelling defects. At the same time, precision is paramount for parametric surfaces, especially when parts are designed to touch (b). Discretizing such surfaces is often undesired, as it may introduce new artifacts (c). Our method can compute winding numbers precisely without surface discretization (d). Truck by Toru Odazawa.
2 Related Work
Winding Numbers
The classical concept of a winding number, originally defined for a closed curve Meister (1769) and easily generalizable to closed surfaces, is one of the topics in complex analysis Needham (1997). Jacobson et al. (2013) generalized winding numbers to polylines and meshes, including the ones with boundary. They also propose an efficient algorithm to compute the generalized winding number that empirically has a time complexity of $O(F^{0.55})$, where $F$ is the number of triangles. Later, winding numbers have been further generalized to, among other representations, point clouds and triangle soups Barill et al. (2018). The method of Barill et al. (2018) can also serve as a very efficient approximation for meshes, achieving $O(\log F)$. Under simple assumptions, on meshes our algorithm has a time complexity between those two, while still being exact (Sec. 7). More importantly, these methods target discretized geometry; applying them to parametric surfaces leads to loss of precision (Fig. 9). Our method can be applied to parametric surfaces directly yielding exact winding numbers. Finally, Feng et al. (2023b) extended the definition to curves on discrete surfaces.
Our method computes a generalized winding number of a surface in any representation that supports ray intersection queries, such as parametric surfaces or meshes. A recent work by Spainhour et al. (2024) offers a practical method of finding a generalized winding number for a rational curve in 2D; their technique cannot be easily extended into 3D. In contrast, we propose a general method for any surfaces or curves supporting ray intersections, including parametric surfaces in 2D and 3D. We are not aware any previous work considering generalized winding number of parametric surfaces.
A very related concept is a solid angle, which for a surface is defined as generalized winding number (without normalization) modulo $4\pi$, so it does not measure the number of turns the surface makes around a point, nor does it change under a flip of surface orientation Feng et al. (2023a). We are focusing on the full winding number that includes the multiplicity and depends on the orientation. The solid angle can be similarly defined and effectively computed for a space curve Knöppel (2018); Chern and Ishida (2024); Binysh and Alexander (2018). Note that in 3D we always compute a winding number of a surface, albeit sometimes not discretized.
Surface Reconstruction from 3D Curve Networks
Inspired by recent progress in 3D curve networks creation Bae et al. (2008); Schmidt et al. (2009); Yu et al. (2021), including industry standard applications in VR Adobe (2023); Google (2023); Gravity Sketch (2023), a series of works addressed the rather difficult problem of surface reconstruction from 3D curve networks. Clean curve networks with precise connectivity are often first broken into oriented closed loops and surfaced Abbasinejad (2011); Zhuang et al. (2013). Each loop can be surfaced via minimization of a variety of energies Welch and Witkin (1994); Bobenko and Schröder (2005); Moreton and Séquin (1992), including Dirichlet energy Schneider and Kobbelt (2001) or Laplacian energy Botsch and Kobbelt (2004). Those two energies are particularly interesting, because their first-order optimality condition is a linear PDE, Laplace and biharmonic equations Bloor and Wilson (1990, 1989). Our method is compatible with this last family of methods, where Boundary Element Methods (BEM) apply. Rough curve networks are often surfaced as a whole Huang et al. (2019); Lin et al. (2022). The desired surfaces can be, depending on the domain, piecewise-smooth Yu et al. (2022), curvature-aligned Bessmeltsev et al. (2012); Pan et al. (2015) or simply smooth Huang et al. (2019); Nealen et al. (2007). These reconstruction methods, producing a mesh or a parametric surface, are all compatible with our winding number computation method.
Ray Casting
Our work is influenced by works on ray casting to voxelize shapes Nooruddin and Turk (2003); Houston et al. (2003). The surfaces they focus on have no true boundary, albeit might be composed of multiple components due to noise, so they perform a inside-outside test; ours may have a boundary, so the definition of insidedness is not binary and is better captured by the notion of generalized winding number.
3 Background
We start by defining a generalized winding number for a smooth [^1] Both the theory and the method we propose apply equally to piecewise smooth surfaces. We only use smoothness for the simplicity of exposition. surface $M$ with a boundary $\partial M$. For brevity, later in the text we write ‘winding number’ instead of ‘generalized winding number’.
For a given surface $M$, the winding number at a point $p\in\mathbb{R}^{3}~{}\setminus~{}M$ is defined as the integral of the differential solid angle $d\Omega(p)$ subtended at $p$ Barill et al. (2018):
[…]
An equivalent formulation of the winding number, which we use at a starting point for our method, is the number of signed intersections of rays with the surface over all directions Jacobson et al. (2013):
[…]
where $S^{2}_{p}$ is a unit sphere centered at $p$, and $\chi(q)$ is the number of signed intersections between ray $\overrightarrow{pq}$ and $M$. By ‘signed intersections’ we mean that an intersection is counted as $+1$ when the ray is intersecting from the back of the surface, i.e., the ray is aligned with the surface normal, and as $-1$ when it is intersecting from the front. So formally,
$\chi(q)=\sum_{j}\mathrm{Sign}(\overrightarrow{pq}\cdot n(r_{j})),$ (2)
where $r_{j}$ are all the intersection points of a ray $\overrightarrow{pq}$ with the surface, and $n(r_{j})$ are the corresponding normals. We refer to $\chi$ as the number of signed intersections.
Figure 3: We define $\chi$ for a ray as the number of signed intersections (Eq. 2) (a). The value of $\chi$ is constant in each region, $A_{1}$ and $A_{2}$, formed by projecting the curve endpoints onto a unit circle. Note that this holds even if the curve is self-intersecting.
4 Algorithm
The input to our algorithm is a manifold surface $M$, with or without boundary $\partial M$. We first describe a generic algorithm, then clarify its implementation for parametric and discrete surfaces in Sec. 5.1.
In this section we outline our basic algorithm that uses multiple rays for a query point. Later (Sec. 5), we build upon it and describe a significantly faster one-shot algorithm, which uses a single ray for multiple query points.
4.1 Computing the Winding Number
Our goal is to compute the winding number at a point $p\in\mathbb{R}^{3}$, i.e., evaluate the integral in Eq. 1 using the boundary $\partial M$. Computing $\chi(q)$ for a single $q$ includes finding all intersections of a ray $\overrightarrow{pq}$ with a potentially non-discretized surface — a rather expensive operation (Sec. 6.2). Therefore, evaluating the integral naïvely is infeasible, as it requires many such ray-surface intersections. Instead, our main observation is the following lemma, a corollary of a more general and technical proposition we prove in the Supplementary materials:
Lemma 4.1
Let $M\subset\mathbb{R}^{3}$ be a smooth surface with a boundary $\partial M$. Let $S^{2}_{p}$ be a unit sphere centered at $p\in\mathbb{R}^{3}$, and let the projection of $\partial M$ onto $S^{2}_{p}$ split the sphere into open regions $A_{i}$. Then for each $i$ there is a constant $\chi_{i}\in\mathbb{Z}$ such that $\chi(q)=\chi_{i}$ almost everywhere in each $A_{i}\ni q$.
Figure 4: Overview of the basic algorithm: Given a surface with a boundary (a), we project the boundary onto a unit sphere around the query point (b, blowup in c). The boundary splits the surface of the sphere into regions (d). We shoot a single ray through each of those regions, and compute the number of signed intersections $\chi_{i}$ of each ray with the surface.
The lemma is illustrated in Fig. 3. Intuitively, this means that $\chi$ is constant within each region, except maybe for a few isolated curves or points. In the figure, we show the 2D example: the surface $M$ becomes a curve, the boundary of $M$ is now a pair of endpoints; their projections onto the unit circle split it into two regions $A_{1}$ and $A_{2}$. The lemma states that even though the number of unsigned intersections may change over each region $A_{i}$ on the sphere, the number of signed intersections is a constant within each region (except for some special points, see below). We denote that constant as $\chi_{i}$. The projection of the boundary itself has measure zero, thus we do not need to define $\chi(q),q\in\textrm{Proj}_{S_{p}^{2}}\partial M$, as these values do not influence the integral. The lemma therefore allows us to rewrite Eq. 1 as a discrete sum:
$w_{M}(p)=\frac{1}{{\color[rgb]{0.0,0.0,0.00}4}\pi}\sum_{i}\textrm{Area}_{S^{2}_{p}}(A_{i})\chi_{i}.$ (3)
The technical detail ‘almost everywhere’ in the formulation of the lemma refers to the points $q$ where the rays $\overrightarrow{pq}$ are tangent to the surface at a contact point (see inset). The set of such points, however, has measure zero, so they do not affect the integral (see Supplementary).
The observation captured in this lemma significantly simplifies our algorithm. Now we can shoot one ray per region $A_{i}$, compute the number of signed intersections $\chi_{i}$, and sum their contributions multiplied by the region area $\textrm{Area}_{S^{2}_{p}}(A_{i})$. Considering that number of regions is typically small, this leads to an efficient algorithm: project the boundary onto the sphere (Fig. 4c), decompose the sphere surface into regions, compute their areas, shoot a single ray through each one and compute $\chi_{i}$ (Fig. 4d).
Figure 5: For a given query point and a unit sphere around it, the projection of the boundary of a surface patch splits the sphere into two ((a), the blue ‘inside’ and the beige ‘outside’) or more regions (b) $A_{i}$ with a constant number of signed intersections $\chi_{i}$. Blue is the ‘front’ of the surface, i.e., the side with the positive normal; the boundary is oriented accordingly.
If the curve network has no boundary, i.e., represents a closed manifold, then the sphere only has one region $A_{1}$, so we compute only one $\chi_{1}$. Such manifolds divide the space into ‘inside’ and ‘outside’ (see Jordan–Brouwer separation theorem Guillemin and Pollack (2010)), so in this case our algorithm is equivalent to a simple inside-outside ray casting test.
For a compact manifold with boundary, the projection of the boundary $\textrm{Proj}_{S^{2}_{p}}\partial M$ on the unit sphere $S^{2}_{p}$ divides the surface of the sphere into two or more regions $A_{i}$. For example, for a surface with a single boundary, $\textrm{Proj}_{S^{2}_{p}}\partial M$ divides the sphere into exactly two regions if the boundary projection has no self-intersections, and more if it has (Fig. 5).
Figure 6: (a) Between two adjacent regions on the sphere, $\chi$ can only change by +1 or -1. In this example, as the ray moves from the bottom ray to the top ray, passing from one region on the sphere to the other, the number of unsigned intersections changes from 2 to 1, and $\chi$ changes from 0 to 1. (b) An example of regions and their $\chi$ values: For each region, $\chi$ is a constant. Furthermore, any two regions $i,j$ adjacent across a segment have $|\chi_{i}-\chi_{j}|=1$. Finally, for a correctly oriented boundary curve, the normals always point towards increasing $\chi$.
5 One-shot algorithm
Performing ray-surface intersection tests for a non-discretized surface is still rather expensive, so we would like to minimize their number. Our key observation enabling an efficient algorithm is the following lemma:
Lemma 5.1
For two regions $i,j$ adjacent across a common edge, $|\chi_{i}-\chi_{j}|=1$.
Indeed, crossing the projection of the boundary $\partial M$ means one of the regions has an additional intersection with $M$ (Fig. 6). This lemma is also a direct corollary of the proposition in the Supplementary. By ‘adjacent along a common edge’ we mean that two regions touching only at a common point do not follow that property. For instance, in Fig. 6, while the $\chi=0$ and $1$ regions are adjacent across edges and thus follow the lemma, the $\chi=1$ and $\chi=-1$ regions only share a common point, so $|\chi_{i}-\chi_{j}|=2$.
If we could know the sign of that difference for every pair of adjacent regions, computing all $\chi_{i}$ would be trivial once we know the value of $\chi$ for at least one region. To determine the sign of the difference $\chi_{ij}\vcentcolon=\chi_{i}-\chi_{j}$, we orient the in-plane normals to the projected boundary curve (Fig. 6) so that they always point towards a higher $\chi$ value. We observe that these normals are changing continuously along the projected curve boundary across the intersections (Fig. 6), and are always making a left turn with respect to the projected curve tangent. Therefore, we define these normals simply as $n_{S^{2}_{p}}(s)\times\tau(s)$.
We first need to compute $\chi_{i}$ for exactly one arbitrary region. We heuristically choose the largest region $i=\mathrm{argmax}_{j}\mathrm{Area}(A_{j})$ and compute $\chi_{i}$ using the ray intersection procedure (Sec. 6.2), using an arbitrary point $q$ within a region. For many points the largest region contains no intersections (e.g., Fig. 6), so most likely a ray intersection test will return quickly.
If we find at least one intersection, we additionally test whether the ray is tangent to the surface at one of those points. If it is, we ignore the ray, because it is a point of contact rather than an intersection, so the number of intersection points for this ray is unreliable, and choose a different random ray through the region. We test this by verifying $|\overrightarrow{pq}\cdot n|>\varepsilon$, where $n$ is the surface normal. For BEM, we use $\varepsilon=10^{-2}$; for parametric geometry, $\varepsilon=10^{-12}$.
Once all the pairwise differences $\chi_{ij}$ are known for all adjacent regions $i,j$, we can represent adjacency between the regions as a directed graph, where each region becomes a vertex, and each pair of adjacent regions $i,j$ gets two edges $i\rightarrow j$ and $j\rightarrow i$, associated with $\chi_{ij}$ and $\chi_{ji}$, respectively. Since it is a connected graph, knowing $\chi_{i}$ for a single ’seed’ vertex $i$ is enough to compute all the other $\chi_{j}$. We therefore use breadth-first search on the graph starting from $i$ and fill in all the other values of $\chi_{j}$, which completes our algorithm. Note that the typical number of regions, and therefore graph vertices, is fewer than 5, so this is essentially instantaneous.
Optimization
We further observe that once we know all the intersections along the ray $\overrightarrow{pq}$, we can immediately compute the winding numbers for all the points along that ray without any more intersection tests. More precisely, once we compute all the ray parameter values $t_{j}$ where it intersects the surface, we know $\chi$ for any point with parameter $t$ along that ray by only counting the intersections with $t_{j}>t$. Thus, for every point along that ray, we only need to perform the decomposition into spherical regions to compute its winding number.
This optimization allows us to very efficiently compute winding numbers in some typical scenarios. For instance, for all points in a planar slice of an object (Fig. 11), where we only need to shoot $\min(W,H)$ rays for a $W$ by $H$ pixel image. Similarly, to compute voxelizations of resolution $N^{3}$, we only need to shoot $N^{2}$ rays (Fig. 12). In those scenarios, this technique allows us to compute winding numbers using fewer than one ray per point on average.
Algorithm 1 One-Shot Generalized Winding Number
1:function Winding Number($p,C$) $\triangleright$ $p\in\mathbb{R}^{3}$, $C$ is a closed curve
2:     $C^{\prime}$ = Project $\partial C$ onto $S^{2}_{p}$
3:     $\{A_{n}\}$ = IdentifyRegions($C^{\prime}$)
4:     $i\leftarrow\textrm{argmax}\;\textrm{Area}(\{A_{n}\})$
5:     $q\leftarrow\textrm{NonTangentRay}(A_{i})$
6:     $\{r_{k},t_{k},n_{k}\}\leftarrow\textrm{Intersect}(\overrightarrow{pq})$triangleright $r_{k}\in\Omega,t_{k}\geq 0,n_{k}\in\mathbb{R}^{3}$
7:     $\chi_{i}\leftarrow\sum_{k}\textrm{Sign}(n_{k}\cdot\overrightarrow{pq})$triangleright Eq. 2
8:     $\{\chi_{ij}\}\leftarrow\textrm{PairwiseDifferences}(\{A_{n}\})$
9:     $\{\chi_{j}\}=\textrm{BFS}(\{\chi_{ij}\},i)$
10:     return $\frac{1}{4\pi}\sum_{j}\chi_{j}\cdot Area(A_{j})$triangleright Eq. 3
11:end function
5.1 Implementation Details
To implement this generic algorithm on a particular geometry representation, we need three main procedures: finding intersections between boundary curves projected onto the unit sphere, ray-surface intersection, and computing areas on the sphere.
Meshes
For meshes, we find ray-mesh intersections using a k-d tree; areas are computed using the spherical polygon area formula as simply $A=2\pi-\sum_{i}(\alpha_{i}-\pi),$ where $\alpha_{i}$ are the interior angles. Finally, we identify regions by computing a surface arrangement on the sphere by first finding intersections between polylines on a sphere, which can be done using a Surface Sweep Algorithm. For a maximum constant number of intersection points, its time complexity is $\mathcal{O}(B\log B)$ Berg et al. (2008), where $B$ is the number of mesh boundary segments. All in all, the asymptotic time complexity of our algorithm is $\mathcal{O}(B\log B+\log F)$ for a single query point. For such scenarios as voxelization, however, the ray intersection is done only one for many query points, so the time complexity for meshes is $\mathcal{O}(B\log B+\frac{1}{\max(W,H)}\log F)$ per point, which for meshes with $B\ll F$ is faster than the state-of-the-art exact winding number computation with the empirical complexity of $\mathcal{O}(F^{0.55})$ Jacobson et al. (2013). Note that while Barill et al. (2018) computes a much faster approximation, our algorithm computes exact winding number.
Parametric Surfaces
For surfaces made out of parametric patches, such as Coons patches, Bézier patches or triangles, or NURBS patches, we can compute the winding number within machine precision without surface discretization as follows. Starting with the problem of ray-parametric patch intersection, let’s take a Bézier triangle as an example, which in essence is a low-degree polynomial $f(u,v)$ with a simplex $u,v>0,u+v<1$ as the parametric domain.
Then for a given ray $O+tR$, where $O,R\in\mathbb{R}^{3},t>0$, we can formulate the following optimization problem that finds the first intersection along the ray:
$\begin{split}&\min_{u,v,t}t\\ &\textrm{s.t.}f(u,v)=O+tR\\ &u,v,t\geq 0\\ &u+v\leq 1.\end{split}$ (4)
As we show in the Supplementary materials, this, as well as the problems of intersecting two parametric curves, such as Bézier, and finding their self-intersections, can be relaxed into a Sum-of-Squares (SOS) formulations Marschner et al. (2021), which then yield convex problems solvable by standard semidefinite programming (SDP) solvers. This relaxation is tight, yielding exact recovery in almost all cases (Sec. 7, Marschner et al. (2021)). All these techniques, as shown in Marschner et al. (2021) are easily extendable to arbitrary NURBS patches.
Finally, for a region bounded by a parametric curve $\gamma(s):[0,L]\rightarrow S_{p}^{2}$ projected onto the unit sphere, using Gauss-Bonnet theorem, we can compute the area inside the curve as $A=2\pi-\int_{0}^{L}k_{g}(s)$, where $k_{g}$ is the geodesic curvature of the boundary. Note that for most parametric boundaries, including Bézier curves, this integral has to be evaluated numerically. Unlike surface discretization, however, this boundary discretization may lead only to minor differences in the final winding number close to the boundary (Sec. 7).
Following this logic, in practice we can discretize the boundary of a parametric surface and compute region decomposition using polyline intersection algorithm, just like in the mesh case. All the results in the paper are computed via in that manner for optimal performance.
Once the arrangement on the sphere is computed, whether for parametric curves or polylines, we find a single point in the largest region Wein et al. (2024).
Speed up
SOS optimization in general can be slow (see Sec. 7). If speed is more important than precision, in practice we can find ray-surface intersections using a nonlinear equation solver. In this case, we can express the ray-surface intersection test as an equation $f(\xi)+(q-p)t=0$, which we solve numerically for $t\geq 0,\xi\in\Omega$, finding all roots. We first can compute an axis-aligned bounding box a parametric patch via SOS Marschner et al. (2021), then intersect it with each ray, giving $t_{min}$ and $t_{max}$; box’s diagonal is $diag$. If the parametric surface has no self-intersections, we then compute all roots by dividing the interval $[t_{\rm{min}},t_{\rm{max}}]$ into a number of subintervals (in our implementation, we use $\max(2,30\frac{t_{\rm{max}}-t_{\rm{min}}}{diag})$ and initializing the solver with the middle of each subinterval. We then consider the roots duplicate, if the distance between them is less than the accuracy of the solver. If surface can have self-intersections, we recommend using SOS optimization for accuracy.
6 Curve Networks
In this section, we demonstrate an application of our algorithm to determining winding number of a curve network, where each patch satisfies a known linear PDE. In the following, we focus on the Laplace’s equation for the simplicity of exposition; our method can be generalized to any PDE where Boundary Element Method is applicable, such as Poisson equation, biharmonic equation, or others.
6.1 Preprocessing
Given an input curve network composed of polylines, we use the method of Zhuang et al. (2013) to detect loops, consisting of input curve segments. We then identify boundary as edges adjacent to only one loop. If the surface is an orientable 2-manifold, we then consistently orient the loops via breadth-first search such that each non-boundary curve segment is traversed once in each direction. This procedure detects if the surface is not a manifold, in which case we compute a winding number as a sum of winding numbers of each manifold patch.
6.2 Ray-Surface Intersection
We need to find a point on the ray $\overrightarrow{pq}$ that satisfies the boundary value problem for a linear PDE for each coordinate $(x,y,z)$.
We formulate the Dirichlet problem for Laplace’s equation for $x:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}$:
$\begin{split}&\Delta x=0\\ &\left.x\right|_{\partial\Omega}=\tilde{x},\end{split}$ (5)
where $\tilde{x}$ is the given curve loop $x$ coordinate. For simplicity, for the examples we set $\Omega=[0,1]^{2}$. Solutions to this problem are known to be minimal surfaces.
Here we follow the Boundary Element Method (BEM) with collocation LaForce (2006). We express the surface as a smooth map $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, where each coordinate of $f(\xi)=(x,y,z)$ is the solution to the corresponding boundary value problem (Eq. 5). In BEM, a representation formula states that the solution of the linear differential equation at a point $\xi\in\Omega$ is expressible as a boundary integral over $\Gamma=\partial\Omega$, which converts our surface into parametric form:
$f(\xi)=\int_{\eta\in\Gamma}G(\xi,\eta)\frac{\partial f(\eta)}{\partial n}-\frac{\partial G(\xi,\eta)}{\partial n}f(\eta)d\Gamma,$ (6)
where $G(\xi,\eta)$ is the fundamental solution of the PDE, and $n$ is the outwards normal of $\Gamma$ (see Supplementary for details).
We discretize the integrals in Eq. 6 using constant boundary elements, and numerically integrate for each polyline segment using trapezoid integration. To find Neumann boundary conditions, i.e., the normal derivative $\left.\frac{\partial f(\xi)}{\partial n}\right|_{\Gamma}$, we use the Dirichlet boundary conditions, which are the curve loop’s coordinates, and solve a dense linear system LaForce (2006) via Cholesky decomposition.
Once both boundary conditions are known, we apply our method using the representation formula in Eq. 6 as a parametric form. We further optimize the process by observing that a minimal surface is contained inside a bounding box of the boundary $\tilde{x}$ due to the maximum principle. We can therefore easily compute the minimum and maximum possible $t$ values $t_{\rm{min}},t_{\rm{max}}$ by intersecting the ray with the axis-aligned bounding box, improving performance.
7 Validation and Results
We implemented our method in C++ with MATLAB subroutines for SOS optimization. We used Eigen Guennebaud et al. (2010), YALMIP Löfberg (2004) as the SOS interface and MOSEK ApS (2024) to solve. We use sphericalpolygon library Li (2020) for areas. We use GNU Scientific Library’s Galassi et al. (1996) implementation of Powell’s Hybrid method as a nonlinear solver when not using SOS.
Throughout the paper, we demonstrate the results of our one-shot winding algorithm on various parametric surfaces (a set of Coons patches in Fig. 1a and 14b, extrusion surfaces in Fig. 2d,14a), meshes (Fig. 1c, 13), and curve networks with minimal surfaces (Fig. 1b, 11). We demonstrate surfaces with boundary, because without boundary our method is a trivial inside-outside test. The boundaries on the curve networks are selected manually. In Fig. 12, we show a typical application of winding numbers, a voxelization of the volume enclosed inside the curve network, computed without discretizing the surfaces. We consider a voxel inside if its center has a winding number $\geq 0.5$. We demonstrate boolean operations in Fig. 15. To compute the ‘set union’, we add the winding numbers; to compute ‘set intersection’, we multiply them pointwise.
SOS accuracy
We validate the accuracy of the SOS solvers, showing all the intersections with parametric surfaces can be found within machine precision. Note that Bézier curves and Bézier triangles were chosen as an example; Marschner et al. (2021); Zhang et al. (2023) demonstrate that all these techniques are easily generalizable to arbitrary NURBS patches.
In each test, we generate the random control points, as well as ray origin and direction, following a uniform $[-1,1]$ distribution. We generated 400 test pairs (half with intersections, half without) for each test. For the ray intersections, we got exact recovery (within machine precision) or correctly identified no intersections for $99.75\%$ of the rays. Exact recovery is identified as a corresponding eigenvalue being larger than a threshold ($10^{-3}$ in our experiments) Marschner et al. (2021). The only ray that did not have exact recovery had an error in $t$ of 0.0064; such rare ray can be simply ignored using the eigenvalue threshold. The projected Bézier curve intersection test has a similar exact recovery rate of 99.75%, same for self-intersections.
Parametric with discretized boundary
We validate the accuracy of our one-shot winding number computation for a discretized boundary and compare it with the accuracy of the hierarchical winding number algorithm (HWN) Jacobson et al. (2013) and Barill et al. (2018) that use a mesh (Fig. 9). As as simple test case, we took a randomly generated Bézier triangle, generated $10^{4}$ query points evenly in an arbitrary slice, and plotted maximum and mean $\ell_{2}$ (RMSE) errors in a logarithmic scale of our method and theirs as functions of the number of boundary samples and mesh vertices, respectively. For the mesh-based methods, we meshed the parametric domain with a regular triangle mesh. The horizontal axis indicates the number of boundary edges for our method and their mesh (bottom), and the corresponding number of faces (top).
Even with a small number of boundary edges, our method is quite precise. Both mesh-based methods, until a certain refinement level, have large maximum errors close to the surface: Each time a query point is on one side of the parametric surface and on the other of the mesh, the winding number will have a wrong sign, yielding a significant error. In contrast, the error of our method, even for roughly discretized boundaries, is significantly smaller, noticeable only around the surface boundary, and stems mainly from approximation of the spherical areas. Using SOS completely eliminates all others sources of error.
As refinement grows, FWN Barill et al. (2018), being an approximation, does not improve. Our method becomes an order of magnitude more precise than the exact mesh-based Jacobson et al. (2013); our worst (maximum) error is roughly the same as their average. Note that to get to a similar maximum error accuracy as our $1900$ boundary edges, mesh-based methods would need roughly $16.3\cdot 10^{6}$ faces, taking 0.9Gb in memory. Our memory footprint is negligible in comparison (kilobytes).
Performance
We evaluate performance of our method on a Intel® Core™ i9-7900X @ 3.30GHz. SOS, while being precise, is slower than our C++ code: finding intersections of projected Bézier curves takes $592\pm 275ms$ via SOS versus $1.2$ms via the polyline intersection code; ray-Bézier triangle intersection takes $581\pm 25$ms via SOS versus $8.85$ms via the nonlinear solver with our parameters. Most of SOS time is spent in YALMIP, not in the solver. We therefore measure all the total performances using the C++ code, parallelized via OpenMP.
We present the results in Table 8 for parametric and minimal surfaces, computing winding numbers of all points in a slice (i.e., the optimization in Sec. 5.1 is on). The performance per point depends mostly on the number of boundary samples; for most of our results a query point takes $\approx 0.2$ms. Dense system solve for BEM is done once for all the query points and takes 10-90ms.
The bottleneck of our algorithm is marked as ’Boundary Processing’ (BP); almost all time is spent on computing finding all intersections between spherical segments. Note that despite the optimal algorithmic complexity of $O(B\log B)$, in our implementation we use a naïve $O(B^{2})$ algorithm which may explain the performance.
For meshes, the accuracy of our method is exactly the same as the ground truth of Jacobson et al. (2013). Performance-wise, our method may bring advantages for meshes where, as discussed in Sec. 5.1, the boundary has few elements, while the mesh has many faces. In Table 8 we present typical performance statistics for such meshes. Note that the approximation algorithm Barill et al. (2018) timing statistics are a few times faster: $0.59,0.55,0.53\mu$s. Our goal, however, is precise winding number computation: as Fig. 9 shows, their approximation is often imprecise.
Input # L # B # Q % B.P. Ours (ms)
Beetle 53 950 196k/308k 93.4 0.157
Hand 132 2150 136k/136k 91.92 0.26
Enterprise 117 1632 250k 99.96 0.26
Spacecraft 20 2761 113k 98.8 1.20
Roll 1 200 250k 95.93 0.10
Car 25 1500 180k 28.92 1.13
Ship 30 1780 262k 20.98 1.84
Gear 2 400 250k 41.66 0.208
Figure 7: Timing statistics for curve networks with minimal surfaces (top half) and parametric surfaces with Coons patches (Car, Ship), and extrusion surfaces (gear, roll). The columns #$\#L,\#B,\#Q$ refer to loops, boundary edges, and query points respectively. Total time in ms per query point is in the last column; time spent boundary processing is %BP.
Input # F # B # Q % B. P. HGWN ($\mu$s) Ours ($\mu$s)
Piggybank 11k 6 600k 96.66 4.41 1.46
Bunny 5k 9 360k 97.64 2.69 2.12
Camel 23k 56 480k 99.01 4.71 8.99
Figure 8: Bottom: statistics for meshes, time in $\mu$s ($10^{-6}s$) per query point. HGWN refers to the hierarchical method of Jacobson et al. (2013).
Curve networks with minimal surfaces
We compare the accuracy of the nonlinear solver approach for a BEM curve network, with the method of Jacobson et al. (2013) on a FEM discretization of the surface using linear elements. As the ground truth surface, we take $z(x,y)=e^{x-1}\sin(y)-e^{x}\cos(y)$; we mesh it and compute its winding numbers. We then reconstruct the surface by FEM using a regular grid on the parametric domain. Then we uniformly sample $342$k points, keeping only those 2368 points with the absolute value of the ground truth winding number $\in(0.4995,0.5005)$; on the rest the error is 0. We can see (Fig. 10) that our solution is as accurate as theirs for just a fraction of the stored data (e.g., 500 vertices vs. 5000 triangles). The discrete increments of the error are because each query point can be classified as either on the correct side of the surface, yielding zero error, and on the wrong side, yielding a winding number error of 1.
Limitations
When a surface has multiple boundaries, such as polygon soups, our assumption of constant number of boundary intersections when projected onto the sphere is no longer true, and the method becomes less efficient.
When not using SOS, the accuracy of our computations relies on the nonlinear equation solver that we use. Errors in root finding, especially missed or spurious roots, can change the value of $\chi$, causing errors. While our analysis (Sec. 7) shows that this rare, for precision applications one should use SOS.
8 Conclusions
We presented a new method for computing a generalized winding number for parametric surfaces, meshes, and curve networks with minimal surfaces. With the use of SOS, our method can compute winding numbers within machine precision without surface discretization. We hope that the future work will explore applications of our method to tracking winding numbers of moving geometry, where discretization becomes even more of an issue.
Figure 9: Logarithmic error plot with respect to ground truth for our method (blue), Jacobson et al. (2013) (HGWN, red) and Barill et al. (2018) (FWN, green). Even when the boundary is discretized, our method computes winding numbers a few orders of magnitude more precisely than Jacobson et al. (2013) or Barill et al. (2018) on a mesh of comparable complexity. Horizontal axis shows the number of faces (top) and boundary vertices (bottom) of a mesh and the number of boundary vertices for our method.
Figure 10: Accuracy of the the winding number computation for curve networks with minimal surfaces computed via BEM. (right). Our method, in general, has similar precision to the mesh-based method (left).
Acknowledgements
We gratefully acknowledge Ivan Puhachov, whose help with figures was invaluable, and to Paul Zhang for his assistance with Sum-of-squares Programming.
Figure 11: Winding numbers computed for each point of a planar slice for the curve networks, surfaced with minimal surfaces, with our one-shot algorithm. Boundaries are highlighted in pink.
Figure 12: The results of our algorithm on closed curve networks, surfaces with minimal surfaces, with our one-shot algorithm. Voxels with a GWN of 1 are occupied and those with a GWN of 0 are left empty.
Figure 13: The results of our one-shot algorithm on meshes. Camel head mesh Sorkine (2005), Vase by virtox.
Figure 14: Additional results on parametric surfaces: a spiral with high winding numbers (a) and a non-manifold boat with Coons patches (b).
Figure 15: Winding numbers enable boolean operations on curve networks. Here ‘set union’ is sum of winding numbers and ‘intersection’ is the pointwise product.
Figure 16: Curve networks can, depending on the context, represent different surfaces, such as (a) parametric patches Bessmeltsev et al. (2012), (b) minimal surfaces Wang and Chern (2021), Weber (2019) (c) other PDEs Nealen et al. (2007), or (d) more complex surfaces Yu et al. (2022).